OpenStudy (anonymous):

Given the function k(x) = x2, compare and contrast how the application of a constant, c, affects the graph. The application of the constant must be discussed in the following manners: k(cx) c • k(x)

OpenStudy (nincompoop):

what do you think it does?

OpenStudy (anonymous):

makes the graph open wider or narrower

OpenStudy (nincompoop):

exactissimo

OpenStudy (nincompoop):

whenever there's a constant that is greater than 1, it means that the function increases faster or sharper, making the graph narrower the same can be said if the constant is greater than -1 the only difference between a positive and negative constant is that it determines whether the parabola opens up or down

OpenStudy (anonymous):

oh i see so what about the second one?

OpenStudy (nincompoop):

I misstyped on the negative constant. instead of greater than -1 I mean lesser than -1

OpenStudy (nincompoop):

for the function f(x) = x^2 f(x) c * x^2 and c * f(x) = x^2 are the same

OpenStudy (anonymous):

i have to go ill be back sorry

OpenStudy (anonymous):

@nincompoop